OpenStudy (anonymous):

the finite region enclosed by the line y=kx, where k is a positive constant, the x axis for0=

OpenStudy (anonymous):



OpenStudy (anonymous):

 The volume of the cone will be \[V=\pi\int_0^h (kx)^2~dx=\frac{k^2\pi}{3}[x^3]_0^h=\frac{k^2h^3\pi}{3}\]